OpenStudy (anonymous):

The sum of two numbers is 44. One number is 2 more than twice the other. Write a system of equations that defines this situation and solve using any method.

OpenStudy (anonymous):

11+33

OpenStudy (anonymous):

the sum of two numbers ( we will use x and y) sum means add so, x + y = 44 one number is 2 more than twice the other , twice being *2 so x = 2y +2 (twice y, and 2 more) substitute that for x in the first equation (2y +2) + y = 44 simplify it, 2y +y =42 ... 3y = 42 divide out by 3. y = 14 put that into the first equation x + 14 = 44 .. solve for x ..x = 30 check if true , 30+ 14 = 44, true 2*14 +2 = 30, true so numbers would be 14, and 30